Ryanmooredis24:

For a function p(x) , as |x| → ∞ then p(x) → 2. Which of the following graphs could be p(x)?

Ryanmooredis24:

1 attachment
Ryanmooredis24:

1 attachment
ThisGirlPretty:

@Vocaloid

Ryanmooredis24:

It says he's offline 🙁

sillybilly123:

\(\begin{align} \lim\limits_{|x| \to \infty} p(x) = 2 \equiv \left\{ \begin{array}{cc} \lim\limits_{x \to \color{red}{-}\infty} p(x) = 2 & \hspace{5mm} x < 0 \\ \lim\limits_{x \to \infty} p(x) = 2 & \hspace{5mm} x \ge 0 \\ \end{array} \right. \end{align}\) right?

sillybilly123:

so you should see 2 at either extreme: \(\pm \infty\)